# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v4

@@ A.1 Unforgeability Proof (Full)
Replace honest shares with uniform: for $j\notin C$, set $\mathbf{s}_{1,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ and $\mathbf{s}_{2,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k}$. Analysis: By Shamir’s $(T,N)$-threshold security, any <T<T shares are information-theoretically independent of the secret. Since $|C|\leq T-1$, the adversary’s view is identical: $\Pr[\mathsf{Win}_{3}]=\Pr[\mathsf{Win}_{2}]$.
@@
Simulate signatures without $\mathbf{s}_{1}$:
@@
    Sample $\mathbf{z}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{v\in R_{q}^{\ell}:\|\mathbf{v}\|_{\infty}<\gamma_{1}-\beta\}$ with appropriate rejection sampling
@@
    Simulate honest parties’ responses $\mathbf{z}_{j}$ by sampling from a high-entropy distribution (justified by Theorem 4.4: conditional min-entropy $\geq n\ell\cdot\log_{2}(2\gamma_{1}/|S|)$ bits) and their masked values $\mathbf{W}_{j},\mathbf{V}_{j}$ as uniform (justified by Lemma 4.9)
Analysis: By Theorem 4.4, each honest party’s response $\mathbf{z}_{j}$ has high conditional min-entropy (statistical, no computational assumptions). The masked commitment and r0-check values $\mathbf{W}_{j},\mathbf{V}_{j}$ are computationally indistinguishable from uniform by Lemma 4.9. Hybrid over $|S\setminus C|$ honest parties: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3}]|\leq n\ell\cdot(1-2\gamma_{1}/(|S|\cdot q))+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$ (the first term from response simulation is negligible for practical $|S|$; the second from mask-based values uses the union bound over all $\binom{N}{2}$ pairwise PRF keys).
@@
Construct M-SIS adversary $\mathcal{B}$:
@@
    $\mathcal{B}$ sets $\rho$ such that $\mathsf{ExpandA}(\rho)=\mathbf{A}$, samples $\mathbf{t}_{1}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k}$
@@
    $\mathcal{B}$ simulates $\mathsf{Game}_{4}$ for $\mathcal{A}$
@@
    If $\mathcal{A}$ outputs valid forgery $(\mu^{*},\sigma^{*})$, apply General Forking Lemma [BN06]
@@
    Rewind to obtain second forgery with same $\mathbf{w}_{1}^{*}$ but different $c_{1}\neq c_{2}$
  6. 6.
    Define $\mathbf{v}=\mathbf{z}_{1}-\mathbf{z}_{2}$ and $\Delta c=c_{1}-c_{2}$
Analysis: Forking succeeds with probability $\geq\Pr[\mathsf{Win}_{4}]^{2}/(q_{H}\cdot|\mathcal{C}|)$. By verification equation:
$\mathsf{HighBits}(\mathbf{A}\mathbf{z}_{i}-c_{i}\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})=\mathbf{w}_{1}\text{ for }i\in\{1,2\}$
Thus $\mathbf{A}\mathbf{v}=\Delta c\cdot\mathbf{t}_{1}\cdot 2^{d}+\boldsymbol{\delta}$ where $\|\boldsymbol{\delta}\|_{\infty}<2\gamma_{2}$. This yields a short vector for the augmented M-SIS instance $\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]$:
$\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]\cdot\begin{bmatrix}\mathbf{v}\\ \Delta c\end{bmatrix}=\boldsymbol{\delta}$
with $\|(\mathbf{v},\Delta c)\|_{\infty}\leq\max(2(\gamma_{1}-\beta),2)<2\gamma_{1}$, solving M-SIS.
A.2 Privacy Proof (Full)
Complete simulator and indistinguishability analysis for Theorem 4.4.
With the Shamir nonce DKG, privacy of the response $\mathbf{z}_{h}$ is statistical (no computational assumptions) and does not require a hybrid argument. The proof simplifies considerably compared to the computational (PRF-based) approach.
@@ A.2 Privacy Proof (Full)
Round 3 Simulation: For each honest $j\in S\setminus C$, sample $\mathbf{z}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ uniformly. This is justified by Theorem 4.4: the real $\mathbf{z}_{h}$ has high conditional min-entropy ($\geq 5\times$ the secret key entropy for $|S|\leq 17$) from the adversary’s perspective, so the simulator’s uniform sampling introduces negligible statistical distance. The $\mathbf{V}_{j}$ values are simulated uniformly by the same PRF-based argument as for $\mathbf{W}_{j}$.
@@
The simulation is statistical for $\mathbf{z}_{j}$ (high min-entropy, negligible statistical distance from uniform) and computationally indistinguishable for $\mathbf{W}_{j}$ (where $\mathbf{W}_{j}=\lambda_{j}\mathbf{w}_{j}+\mathbf{m}_{j}^{(w)}$) and $\mathbf{V}_{j}$ (PRF security). The total distinguishing advantage is $n\ell\cdot(1-2\gamma_{1}/(|S|\cdot q))+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$, where the first term (negligible for practical $|S|$) arises from the response simulation and the second from mask-based values. In particular, no PRF assumption is needed for response privacy.
@@ Appendix B Distributed Key Generation Protocols
Our protocol uses two structurally similar but functionally distinct DKG protocols: a key DKG (executed once during setup) and a nonce DKG (executed before each signing session, typically preprocessed offline). Both follow the Feldman VSS paradigm but differ in their input distributions and security requirements.
@@ B.2 Nonce DKG (Per-Session Preprocessing)
The protocol is specified formally in Algorithm 1 of the main paper (Section 3.4). Each party $i\in S$ generates a degree-(T1)(T\!-\!1) polynomial $\hat{f}_{i}(X)$ whose constant term $\hat{\mathbf{y}}_{i}^{(0)}$ is sampled from the reduced nonce range $\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$ and whose higher-degree coefficients are sampled uniformly from $R_{q}^{\ell}$. The resulting aggregated nonce $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}^{(0)}$ follows the Irwin-Hall distribution analyzed in Theorem 4.11, while each honest party’s individual share $\mathbf{y}_{h}=\sum_{i\in S}\hat{f}_{i}(h)$ has high conditional min-entropy from the adversary’s perspective (Theorem 4.4).
@@ B.3 Security Properties
  • Robustness: With Feldman commitments [Feldman87], malicious shares can be detected and attributed.
The nonce DKG additionally provides statistical privacy (no computational assumptions) of individual shares (Theorem 4.4), which is the key enabler for eliminating the two-honest requirement for response privacy in coordinator-based profiles (P1, P3+; see Remark 3.7). The key DKG does not require this property, since key shares are never revealed during signing.
@@ Appendix D Numerical Verification of Rényi Bounds
For signing set size $|S|=n$, each party samples nonces from range $N=2\lfloor\gamma_{1}/n\rfloor+1$. For readability, we use the approximation $N\approx 2\gamma_{1}/n$ (the difference is at most $1/N<2\times 10^{-5}$ for all relevant $n$, and the exact discrete range is used in the implementation). Using Lemma 4.2 of [Raccoon2024] with $\alpha=2$ (where $\alpha(\alpha-1)/2=1$) and shift $c=\beta$:
$R_{2}^{\epsilon}\leq 1+\frac{n^{2}\cdot\beta^{2}}{N^{2}}=1+\frac{n^{2}\cdot\beta^{2}\cdot n^{2}}{4\gamma_{1}^{2}}=1+\frac{n^{4}\cdot\beta^{2}}{4\gamma_{1}^{2}}$
$|S|$ $N=2\gamma_{1}/|S|$ $R_{2}^{\epsilon}-1$ $\epsilon$ (tail)
@@
17 61681 $2.9\times 10^{-3}$ $10^{-30}$
25 41943 $1.4\times 10^{-2}$ $10^{-40}$
33 31775 $4.1\times 10^{-2}$ $10^{-49}$
Table 12: Per-coordinate Rényi divergence bounds for ML-DSA-65. All values computed using proven Lemma 4.2 of [Raccoon2024]. Full-vector: $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$.
The per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}$ remains below $1.05$ for $|S|\leq 33$. The full-vector divergence $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$ ranges from $\approx 2^{5.4}$ ($|S|=17$) to $\approx 2^{75}$ ($|S|=33$); see Theorem 4.11 for the impact on the security bound.

### v5

@@ A.1 Unforgeability Proof (Full)
We argue that the adversary’s view of the honest parties’ key shares $\{\mathbf{s}_{1,j}\}_{j\notin C}$ is statistically independent of the secret key $\mathbf{s}_{1}$, by Shamir’s $(T,N)$-threshold perfect secrecy: any T1T-1 evaluations are jointly consistent with every possible value of the secret (the marginal distribution of each honest share at a non-zero evaluation point is $R_{q}^{\ell}$-uniform, but the joint distribution of multiple honest shares is not i.i.d.—they are correlated as evaluations of the same degree-(T1)(T\!-\!1) polynomial; the correct statement is that the joint view reveals zero information about $\mathbf{s}_{1}$). Important: this is a conceptual argument about the adversary’s view—the signing oracle continues using the actual short key shares (replacing them with uniform $R_{q}^{\ell}$ values would cause the $\|\mathbf{z}\|_{\infty}$-bound and r0-checks to fail with overwhelming probability, invalidating signatures). Since $|C|\leq T-1$, the adversary’s view of honest shares is statistically independent of the secret: $\Pr[\mathsf{Win}_{3}]=\Pr[\mathsf{Win}_{2}]$. (In Profile P2, individual per-party responses $\mathbf{z}_{j}$ are contributed only as SPDZ input shares, so SPDZ input-privacy ensures the adversary observes only the aggregate $\mathbf{z}$, not individual $\mathbf{z}_{j}$ values; the same Shamir argument applies.)
Game 3.5 (Uniform Nonces).
Replace Irwin-Hall nonce distribution with uniform over $\{-\gamma_{1},\ldots,\gamma_{1}\}^{n\ell}$. In Game 3, the aggregate nonce $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ follows the Irwin-Hall distribution. We replace this with $\mathbf{y}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1},\ldots,\gamma_{1}\}^{n\ell}$ uniform. Analysis: The $q_{s}$ signing sessions are independent, so the joint Rényi divergence is $(R_{2}^{\mathsf{vec,shift}})^{q_{s}}$ (direct shift-invariance, Theorem Q.25). Applying the smooth Rényi divergence probability transfer lemma (Lemma Q.13) to the joint distribution:
$\Pr[\mathsf{Win}_{3}]\leq\bigl(R_{2}^{\mathsf{vec,shift}}\bigr)^{q_{s}/2}\cdot\sqrt{\Pr[\mathsf{Win}_{3.5}]}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
where $R_{2}^{\mathsf{vec,shift}}=(1+\chi^{2}_{\mathsf{direct}})^{n\ell}$ is the full-vector per-session Rényi divergence from the direct shift-invariance analysis, with $\chi^{2}_{\mathsf{direct}}\approx 6.62\times 10^{-6}$ per coordinate for $|S|=17$ (Corollary Q.24), giving $R_{2}^{\mathsf{vec,shift}}\approx 1.0085$, and $\epsilon_{\mathsf{IH}}<10^{-30}$ (Theorem 4.17). The total security loss is $\approx 6.1\times 10^{-3}\cdot q_{s}$ bits, giving a proven security bound of $\geq 96-0.006\cdot q_{s}$ bits for $|S|\leq 17$ (where the 96-bit baseline is the Cauchy-Schwarz halving of the 192-bit NIST Level 3 baseline). The conservative analytical bound ($R_{2}^{\mathsf{vec}}\approx 2^{5.4}$, Remark 4.5) is vacuous at $q_{s}\gtrsim 36$.
@@
Starting from uniform nonces (Game 3.5), simulate signatures without $\mathbf{s}_{1}$:
@@
    Sample $\mathbf{z}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{v\in R_{q}^{\ell}:\|\mathbf{v}\|_{\infty}<\gamma_{1}-\beta\}$ with rejection sampling
@@
    Simulate masked values $\mathbf{W}_{j},\mathbf{V}_{j}$ as computationally uniform (Lemma 4.15, replacing honest-honest PRF terms with uniform)
Clarification: In this EUF-CMA game, the adversary observes only the aggregate public signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$; individual per-party responses $\mathbf{z}_{j}$ are processed inside the TEE/combiner and never reach the adversary. Accordingly, Steps 1–3 produce the adversary-visible $\mathbf{z}$ without using $\mathbf{s}_{1}$, and Step 4 handles the masked intermediate values that the adversary also cannot observe directly. The full per-party simulation (including $\mathbf{z}_{j}=\mathbf{y}_{j}+c\mathbf{s}_{1,j}$ from DKG values) is deferred to the UC/privacy simulator in Appendix A.2. Analysis: The masked commitment and r0-check values $\mathbf{W}_{j},\mathbf{V}_{j}$ are computationally indistinguishable from uniform by Lemma 4.15. By PRF security: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3.5}]|\leq\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
@@
Construct SelfTargetMSIS adversary $\mathcal{B}$ [KLS18]:
@@
    $\mathcal{B}$ sets $\rho$ such that $\mathsf{ExpandA}(\rho)=\mathbf{A}$, samples $\mathbf{t}_{1}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k}$ (uniformly random; valid since $\mathsf{Game}_{4}$ simulation does not require knowing $\mathbf{s}_{1}$ or $\mathbf{s}_{2}$; $\mathbf{t}_{1}$ is indistinguishable from a real key by M-LWE hardness [KLS18])
@@
    $\mathcal{B}$ picks a random index $i^{*}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}[q_{H}]$ as the target ROM query
@@
    $\mathcal{B}$ simulates $\mathsf{Game}_{4}$ for $\mathcal{A}$, programming ROM responses uniformly
@@
    When $\mathcal{A}$ outputs forgery $(\mu^{*},\mathbf{z}^{*},c^{*},\mathbf{h}^{*})$: if $c^{*}$ equals the $i^{*}$-th ROM response, extract; otherwise abort
Analysis: $\mathcal{B}$ succeeds when its guess $i^{*}$ is correct, which occurs with probability $1/q_{H}$, so $\Pr[\mathcal{B}\ \text{succeeds}]\geq\Pr[\mathsf{Win}_{4}]/q_{H}$. When successful, the forgery verification gives:
$\mathsf{HighBits}(\mathbf{A}\mathbf{z}^{*}-c^{*}\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})=\mathbf{w}_{1}^{*}$
Rearranging yields the SelfTargetMSIS solution for $\mathbf{A}\mid-\mathbf{t}_{1}\cdot 2^{d}]$ (see Remark 2.3):
$\mathbf{A}\mid-\mathbf{t}_{1}\cdot 2^{d}]\cdot\begin{bmatrix}\mathbf{z}^{*}\\ c^{*}\end{bmatrix}=2\gamma_{2}\mathbf{w}_{1}^{*}+\mathbf{r}_{0}^{*}$
where $\mathbf{w}_{1}^{*}=\mathsf{HighBits}(\mathbf{A}\mathbf{z}^{*}-c^{*}\mathbf{t}_{1}\cdot 2^{d},\,2\gamma_{2})$ is the known verification target and $\|\mathbf{r}_{0}^{*}\|_{\infty}\leq\gamma_{2}$ is the $\mathsf{LowBits}$ residual. The combined solution $(\mathbf{z}^{*});(c^{*})]\in R_{q}^{\ell+1}$ satisfies $\|\mathbf{z}^{*}\|_{\infty}<\gamma_{1}-\beta$ (z-bound check) and $\|c^{*}\|_{\infty}=1$ (challenge coefficients in $\{-1,0,+1\}$); the module $\ell^{\infty}$-norm is $\max(\|\mathbf{z}^{*}\|_{\infty},\|c^{*}\|_{\infty})<\gamma_{1}$, matching the SelfTargetMSIS norm parameter from Remark 2.3 and [KLS18]. Thus $\Pr[\mathsf{Win}_{4}]\leq q_{H}\cdot\epsilon_{\mathsf{M\text{-}SIS}}$.
A.2 Privacy Proof Details
We show that the simulator produces an indistinguishable view, applying the min-entropy bound established in Theorem 4.7 (Section 4).
With the Shamir nonce DKG, privacy of the nonce share $\mathbf{y}_{h}$ is statistical (no computational assumptions). The simulator computes responses from simulated DKG values, achieving perfect indistinguishability for the response component.
@@ A.2 Privacy Proof Details
Round 3 Simulation: The simulator must produce per-party responses $\mathbf{z}_{j}$ consistent with both the simulated DKG values and the public signature constraint $\sum_{j\in S}\lambda_{j}\mathbf{z}_{j}=\mathbf{z}$ (enforced by the public $\sigma$). We handle the two cases separately.
Case $|S\setminus C|>1$: Let $\mathbf{z}_{j}^{*}=\mathbf{y}_{j}^{*}+c\cdot\mathbf{s}_{1,j}$ denote corrupted party $j$’s response, where $\mathbf{y}_{j}^{*}$ is $j$’s simulated DKG evaluation (known to the simulator from the corrupted view). The simulator chooses $\mathbf{z}_{j}=\mathbf{y}_{j}+c\cdot\mathbf{s}_{1,j}$ (from simulated DKG values) for all honest $j\in S\setminus C$ except a designated party $h_{0}$, then sets
$\mathbf{z}_{h_{0}}=\frac{\mathbf{z}-\sum_{j\in C}\lambda_{j}\mathbf{z}_{j}^{*}-\sum_{j\in(S\setminus C)\setminus\{h_{0}\}}\lambda_{j}\mathbf{z}_{j}}{\lambda_{h_{0}}}$
(The division by $\lambda_{h_{0}}$ is well-defined: since all evaluation points lie in $\{1,\ldots,N\}\subset\mathbb{Z}_{q}^{*}$, the Lagrange coefficient $\lambda_{h_{0}}=\prod_{j\in S,j\neq h_{0}}j(j-h_{0})^{-1}\bmod q$ is a product of nonzero field elements and hence nonzero mod $q$, hence invertible in $\mathbb{Z}_{q}$.) This $\mathbf{z}_{h_{0}}$ matches exactly what the real protocol would produce (given the DKG polynomials), so SD =0=0.
Case $|S\setminus C|=1$: The single honest response $\mathbf{z}_{h}=(\mathbf{z}-\sum_{j\in C}\lambda_{j}\mathbf{z}_{j}^{*})/\lambda_{h}$ is uniquely determined by $\sigma$ and the corrupted views. The simulator outputs this value directly. Privacy of $\mathbf{s}_{1,h}$ (key privacy, computational) is maintained because the adversary knows $\mathbf{z}_{h}$ but not the decomposition $\mathbf{z}_{h}=\mathbf{y}_{h}+c\mathbf{s}_{1,h}$: recovering $\mathbf{s}_{1,h}$ from $\mathbf{z}_{h}$ requires solving a bounded-distance decoding instance given $\mathbf{y}_{h}$’s high conditional min-entropy (Theorem 4.7).
The $\mathbf{V}_{j}$ values are computed as $\mathbf{V}_{j}=\lambda_{j}c\mathbf{s}_{2,j}+\mathbf{m}_{j}^{(s2)}$ using simulated key shares and PRF masks.
@@
In both cases, the simulated $\mathbf{z}_{j}$ follows the same distribution as the real protocol (SD =0=0 conditioned on $\sigma$). Masked values $\mathbf{W}_{j}$ and $\mathbf{V}_{j}$ are computationally indistinguishable from uniform (PRF security, Lemma 4.15); no PRF assumption is needed for nonce share privacy ($\mathbf{y}_{h}$’s statistical privacy is independent). The total distinguishing advantage is $\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
@@ Appendix B Distributed Key Generation Protocols
Our protocol uses two structurally similar but functionally distinct DKG protocols: a key DKG (executed once during setup) and a nonce DKG (executed before each signing session, typically preprocessed offline). The key DKG uses Feldman-style commitments for verifiability; the nonce DKG omits VSS verification (correctness is instead ensured at the signing stage via the z-bound and r0-check). Both share the same polynomial evaluation structure but differ in their input distributions and security requirements.
@@ B.2 Nonce DKG (Per-Session Preprocessing)
The protocol is specified formally in Algorithm 1 of the main paper (Section 3.4). Each party $i\in S$ generates a degree-(T1)(T\!-\!1) polynomial $\hat{f}_{i}(X)$ whose constant term $\hat{\mathbf{y}}_{i}$ is sampled from the reduced nonce range $\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$ and whose higher-degree coefficients are sampled uniformly from $R_{q}^{\ell}$. The resulting aggregated nonce $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ follows the Irwin-Hall distribution analyzed in Theorem 4.17, while each honest party’s individual share $\mathbf{y}_{h}=\sum_{i\in S}\hat{f}_{i}(h)$ has high conditional min-entropy from the adversary’s perspective (Theorem 4.7).
@@ B.3 Security Properties
  • Robustness: The key DKG uses Feldman commitments [Feldman87] to detect and attribute malicious shares. The nonce DKG omits VSS verification; malicious behavior is instead detected at the signing stage (z-bound or r0-check failure triggers the blame protocol), which suffices for security-with-abort.
The nonce DKG additionally provides statistical privacy (no computational assumptions) of individual shares (Theorem 4.7), which is the key enabler for eliminating the two-honest requirement (enabling $|S|=T$ signing sets with nonce share privacy) in coordinator-based profiles (P1, P3+; see Remark 3.9). The key DKG does not require this property, since key shares are never revealed during signing.
@@ C.1 Share Refresh Protocol
Forward Secrecy for Past Signatures.
Signatures from epoch $k$ remain secure after the epoch-(k+1)(k\!+\!1) refresh. A signature $\sigma$ from epoch $k$ reveals $\mathbf{z}_{j}^{(k)}=\mathbf{y}_{j}^{(k)}+c\cdot\mathbf{s}_{1,j}^{(k)}$. Since $\mathbf{y}_{j}^{(k)}$ is a one-time nonce share (deleted after use), and the refreshed shares $\mathbf{s}_{1,j}^{(k+1)}=\mathbf{s}_{1,j}^{(k)}+\text{(zero-sharing term)}$ do not reveal $\mathbf{s}_{1,j}^{(k)}$ by Shamir’s threshold property (the zero-sharing randomizes shares while preserving the secret), past key shares cannot be reconstructed from refreshed shares alone. Key privacy for past signatures therefore holds across epochs under the same M-LWE hardness assumption as the signing scheme.
@@ Appendix D Numerical Verification of Rényi Bounds
Let $m=|S|$ denote the signing set size (distinct from the ring degree $n=256$) and $B=2\lfloor\gamma_{1}/m\rfloor$ the per-share full range (from $\lfloor\gamma_{1}/m\rfloor$ to $\lfloor\gamma_{1}/m\rfloor$, so the total width is $2\lfloor\gamma_{1}/m\rfloor$). Note that in Definition Q.7, the symbol $B$ denotes the half-range (each $X_{i}\in\{-B,\ldots,B\}$); to apply Raccoon [Raccoon2024] Lemma 4.2, $B$ here is the full range, equal to $2\times$ that half-range. For readability, we use the approximation $B\approx 2\gamma_{1}/m$ (the rounding error is at most $2/B<4\times 10^{-5}$ for all relevant $m$, and the exact discrete range is used in the implementation). Using Lemma 4.2 of [Raccoon2024] with $\alpha=2$ and shift $\delta=\beta$:
$R_{2}^{\epsilon}$ $\leq 1+\frac{m^{2}\cdot\beta^{2}}{B^{2}}\approx 1+\frac{m^{2}\cdot\beta^{2}}{(2\gamma_{1}/m)^{2}}$
$=1+\frac{m^{4}\cdot\beta^{2}}{4\gamma_{1}^{2}}=1+\frac{|S|^{4}\cdot\beta^{2}}{4\gamma_{1}^{2}}$
$|S|$ $B\approx 2\gamma_{1}/|S|$ $R_{2}^{\epsilon}-1$ $\epsilon$ (tail)
@@
17 61680 $2.9\times 10^{-3}$ $10^{-30}$
25 41942 $1.4\times 10^{-2}$ $10^{-40}$
33 31774 $4.1\times 10^{-2}$ $10^{-49}$
Table 12: Per-coordinate Rényi divergence bounds for ML-DSA-65. $B=2\lfloor\gamma_{1}/|S|\rfloor$ is the exact per-share full range; displayed values are $2\lfloor\gamma_{1}/|S|\rfloor$. All bounds computed using Lemma 4.2 of [Raccoon2024]. Full-vector: $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$ ($n=256$, $\ell=5$).
The per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}$ remains below $1.05$ for $|S|\leq 33$. The full-vector divergence $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$ ranges from $\approx 2^{5.4}$ ($|S|=17$) to $\approx 2^{75}$ ($|S|=33$); see Theorem 4.17 for the impact on the security bound. The direct analysis (Corollary Q.24) gives <0.013<0.013 bits for all $|S|\leq 33$.